OpenStudy (anonymous):

2(1.05)^(x)+3=10 solve using log

OpenStudy (anonymous):

So it's \((2 \times 1.05^x) + 3 = 10\)?

OpenStudy (anonymous):

\[2(1.05)^x+3=10\]

OpenStudy (anonymous):

Ok, same thing. :) So first you want to get the exponent by itself so you can take the log. I assume you know how to do that.

OpenStudy (anonymous):

I am not sure lol xlog5.1=1??

OpenStudy (anonymous):

I meant do this: \[2(1.05)^x = 7\]\[(1.05)^x = 3.5\]

OpenStudy (anonymous):

dude logs are not my friends

OpenStudy (anonymous):

sry i took so long to respond work gets busy then its dead

OpenStudy (anonymous):

It's ok. Logs are nobody's friends. But they can be helpful. Can you solve it from what I gave you or do you need more?

OpenStudy (anonymous):

yeah one more step lol

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

\[x = log_{1.05} 3.5\]

OpenStudy (anonymous):

okay but how do you solve that with a ti 83 plus

OpenStudy (anonymous):

Use the change of base formula. \[\log_b n = \frac{\log_a n}{\log_a b}\](For any a \(\ne\) 1 and b \(\ne\) 1)

OpenStudy (anonymous):

wow thank you just remembered my teacher showing us that formula will not forget that again